# On the Fundamental Impossibility of Hallucination Control in Large Language Models

### v4

@@ 5 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, we will address some philosophical aspects of the problem, its generality, scope and consequences for the AI development.
@@ 5.1 Knowledge and Semantic Information
Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction effort in a given context. We shall assume finiteness of knowledge representation inspired by the embedding space properties.
That way we adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems. This differs from the probabilistic credence perspective, treating knowledge as justified true beliefs. One way to justify the pragmatic or mechanistic approach focusing on structural transformations rather than belief updates, we can refer to the actual way LLMs process information. In other words, we shall not claim any ontological truth about knowledge, but model what LLMs can operationally achieve.
Now, let us present formal definitions of that ideas.
@@
It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings. For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements.
Elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales. A pragmatist view is to interpret element of knowledge as activation patterns (ideas) encoding anything from individual token embeddings to complex sentences or compositional structures. There is an important assumption that we make that requires addressing. We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space. That limits applicability of our conclusions to inference mechanisms recombining existing patterns. As we will see below, that is not as limiting an assumption as it may seem at first glance.
@@
For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge. For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference.
To describe the inference or reasoning process, we introduce the concept of semantic information measure that registers information processed within available computational budget.
@@
The semantic information measure is a bounded mapping $\mu_{C}:\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ parameterized by computational budget $C$ satisfying the following axioms:
$\mu_{C}(\emptyset)=0\text{ and }\mu_{C}(A)>0\text{ for }A\neq\emptyset,$ (null set and non-negativity)
$\mu_{C}(A)\leq\mu_{C}(B)\text{ if }A\subseteq B,$ (monotonicity)
$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)\text{ with equality for }A\perp_{C}B,$ (subadditivity)
$\mu_{C}(A)\leq\mu_{D}(A)\text{ if }C<D,$ (insight monotonicity)
$\sup\{\mu_{C}(A)\mid A\subseteq\mathcal{K}\}<\infty,$ (boundedness)
where $A\perp_{C}B$ means that no element of $A$ can derive elements of $B$ within computational budget $C$, and vice versa.
The idea of computational independence allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way. We need that distinction to deal with the mechanics of reasoning process. Also, as we will see in the following sections, it plays an important role in our study of the semantic information conservation and properties of bounded reasoning.
Definition 4 (Computational Independence).
Let $\mathcal{F}_{C}$ denote the set of all computational transformations with complexity bounded by $C$. Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
We limit our investigations to those cases for which the following model applies. Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query ($A\cap q=\emptyset$) has zero information. If that is not the case and there is information within our reach that reduces uncertainty, we may hope for some content useful for reasoning. To make that possible, we need to allow for information ordering. First, we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa. Second, with redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation. Finally, we say that with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
5.1.1 Justification for the Axioms of the Semantic Information Measure
The choice of conditional additivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas. As we will see, we need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns. The concept of Computational Independence is the formal translation of this informational separation into a computational context. It formally states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
@@ 5.1.1 Justification for the Axioms of the Semantic Information Measure
This axiomatic choice is most applicable to the idealized setting of Theorem 6. Later, in Theorem 8 and 9, we relax that model by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
5.1.2 Justification for the Axiom of Unlimited Reasoning with Emergence Operator
That axiom is essential for the proof of idempotence of unlimited reasoning in Theorem 3 Theorem 4. These are critical philosophical and mathematical results necessary for our study. With them we argue that the information gain observed in practical, budget-constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo.
@@ 5.1.3 Limits of Knowledge and Reasoning
For any $A\subseteq\mathcal{P}(\mathcal{K})$ there exists
$\mu_{\infty}(A)=\lim_{C\to\infty}\mu_{C}(A).$ (2)
@@
Since $\mu_{C}(A)\leq\mu_{D}(A)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space. ∎
@@
Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ if $C<D$ and such that for any $g,h\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$. Emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (3)
@@
$\mathcal{E}_{\infty}(A)=A\cup\{B\in\mathcal{K}\mid B=f(A)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (4)
@@
Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest. When $A\subseteq\mathcal{E}_{C}(A)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible. That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
When $\mathcal{E}_{C}(A)=A\cup B$, we say that $B$ emerges from reasoning based on $A$. Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$. In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information as expected. Whenever $A\subseteq A\cup B$, we have
$\mu_{C}(A)\leq\mu_{C}(A\cup B)=\mu_{C}(\mathcal{E}_{C}(A))\leq\mu_{C}(A)+\mu_{C}(B).$ (6)
The emergence operator is monotonic, which means the reasoning process it describes is also monotonic. Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{C}(B)$. Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A)\subseteq\mathcal{E}_{D}(A)$. By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
@@
For any $A\subseteq\mathcal{P}(\mathcal{K})$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A)=\mathcal{E}_{C}(A^{*}_{C}).$ (7)
@@
$\mu_{\infty}(A)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A)|.$ (8)
@@
$\mu_{\infty}(\emptyset)=\log|\mathcal{K}|\cdot|\bigcup_{n}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
If $A\subseteq B$, then $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(B)$ implies $\mu_{\infty}(A)\leq\mu_{\infty}(B)$. When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa). Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$. That means $\mu_{\infty}(A\cup B)=\mu_{\infty}(A)+\mu_{\infty}(B)$ for computationally independent sets. That shows the proposed formula satisfies all required properties of semantic information measure.
@@
We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent, $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y)$ and $Y\subseteq\mathcal{E}_{\infty}(X)$.
Notice that equivalence of knowledge does not mean $A$ is a fixed point if emergence, i.e., $A=\mathcal{E}_{\infty}(A)$ does not mean $A\equiv\mathcal{E}_{\infty}(A)$. If we have $A=\mathcal{E}_{\infty}(A)$, then nothing more can be derived from $A$ by any reasoning rule from $A$. That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A)$. Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)$. Then $\mathcal{E}_{\infty}(A)=\{x,y\}\supset\{x\}=A$, but $A\equiv\mathcal{E}_{\infty}(A)$.
@@
For any $A\subseteq\mathcal{K}$ we have:
$\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A)).$ (9)
@@
By definition $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$. To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A)=g(A\cup\bigcup f_{i}(A))$ for $g,f_{i}\in\mathcal{F}_{\infty}$. But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any conclusion $g(\mathcal{E}_{\infty}(A))$ derived from $\mathcal{E}_{\infty}(A)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$. Therefore, $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))\subseteq\mathcal{E}_{\infty}(A)$ as required. ∎
@@
For any $A\subseteq\mathcal{K}$, if $\mu_{\infty}$ is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X)=\mu_{\infty}(Y)$ whenever $X\equiv Y$, then:
$\mu_{\infty}(\mathcal{E}_{\infty}(A))=\mu_{\infty}(A).$ (10)
@@
We must show equivalence of knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$. But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ $\mathcal{E}_{\infty}(A)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))=\mathcal{E}_{\infty}(A)$ as requested. Since $A\equiv\mathcal{E}_{\infty}(A)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge, we conclude that $\mathcal{E}_{\infty}(A)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A))$. ∎
The Uncertainty Resolution Paradox and its Resolution
We address the general settings in which any subset of knowledge encoded in LLM can be redundant, so
$\mu_{C}(A\cup B)\leq\mu_{C}(A)+\mu_{C}(B).$
But when LLM combines knowledge elements to fabricate facts or produce emergent insights, it seems that
$\mu_{C}(A\cup B)>\mu_{C}(A)+\mu_{C}(B),$
so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume otherwise?

### v5

@@ 4 Theoretical Framework
This section develops the mathematical foundations necessary for proving the impossibility theorem. It formalizes key concepts including knowledge representation, semantic information, and the auction-theoretic model of LLM inference. Also, we address some philosophical aspects of the reasoning problem, relevant to our study, its generality, scope and consequences for the AI development.
@@ 4.1 Knowledge and Semantic Information
Let us define knowledge as organized information that LLMs can encode and utilize for reasoning, and aligning with epistemological concepts while maintaining mathematical rigor. In other words, knowledge, much like energy (with low entropy) in physics, is a quantity we can use to perform work of reasoning. Information is a (measure of) potential for uncertainty reduction in a given context. We assume finiteness of knowledge representation to refer directly to the properties of the embedding space (knowledge representation space) in LLMs.
We adopt a pragmatist perspective on knowledge, treating it as organized information patterns accessible to computational systems. That differs from the probabilistic credence perspective, natural in Bayesian learning, treating knowledge as justified true beliefs updated based on evidence. That is a rather bold ontological claim, but as we will see, one that allows us to make progress in understanding what LLMs can operationally achieve.
Now, let us present formal definitions of these ideas.
@@
It is natural to think of the knowledge space as the space of embeddings with additional measure and mappings. For model $\mathcal{M}$ with parameters trained on data, $\mathcal{K}_{M}\subset\mathcal{K}$ represents the knowledge subset encoded in the model’s parameters. The distance function $d_{\mathcal{K}}$ measures differences between knowledge elements. The elements of the knowledge space represent atomic facts, relational structures, and transformation rules at many scales. A pragmatist view is to interpret elements of knowledge as activation patterns (ideas) encoding anything from individual tokens of texts to complex sentences or compositional structures.
There is an important (ontological) assumption here that calls for our attention. We purposefully assume $\mathcal{K}$ is closed under transformations available in computable function space. At first glance, that assumption may look as limiting the applicability of our conclusions to combinatorial inference mechanisms, merely recombining subsets of knowledge. That intuition is misleading. In the following sections we introduce the notion of emergence operator to describe reasoning process capable of making novel insights in that setting.
@@
For parameterization $W\in\mathcal{W}$, $\Phi(W)=\mathcal{K}_{M}$ represents encoded knowledge. For activation pattern $a_{i}\in\mathcal{A}$ with parameters $W_{i}$, $\Psi(a_{i})\subset\Phi(W_{i})$ represents knowledge accessed during specific inference. Knowledge representation models how LLMs encode information in their parameters and retrieve it from them.
We need to introduce an idea of computational independence that allows for distinguishing between facts derivable from other facts and facts that cannot be derived that way. We need that distinction to deal with the mechanics of reasoning process.
Definition 3 (Computational Independence).
Let
$\mathcal{F}_{C}=\{f:\mathcal{P}(\mathcal{K})\to\mathcal{P}(\mathcal{K})\mid\text{cost}(f)\leq C\}$ (1)
denote the set of all computational transformations with computational cost bounded by $C$. Knowledge sets $A,B\subseteq\mathcal{K}$ are computationally independent with respect to computational budget $C$, denoted $A\perp_{C}B$, if and only if for all $f\in\mathcal{F}_{C}$:
To gain some intuitions about $\mathcal{F}_{C}$, let us consider several examples. One is the set of all derivation rules in first-order logic, like modus ponens or universal instantiation, with at most $C$ inference steps. Then, give an axiom set $A\subset\mathcal{K}$, we can produce new subset of knowledge with at most $C$ applications of the admitted inference rules. For example, given $A=\{P,P\to Q\}$ and $B=\{R,R\to S\}$ and the budget $C=2$ for two inference steps, we can derive $\{P,P\to Q,Q\}$ $\{R,R\to S,S\}$ with any $f\in\mathcal{F}_{2}$. Then $A\perp_{2}B$ because $f(A)\cap B=\emptyset$ and $A\cap f(B)=\emptyset$, there are no shared elements of knowledge accessible in 2-step derivations.
The computational cost (or complexity) measure $\text{cost}(f)$ could therefore be: Kolmogorov complexity giving the shortest program computing $f$, longest path in computation graph (or length of the chain of thoughts CoT), number of inference rules applied, or number of QKV matrix multiplications in transformers. Then we can define a context-relevant subset of knowledge that can contribute to reasoning within a given context (of a query) within budget some computational budget.
Definition 4 (Context-Relevant Knowledge).
The context-relevant knowledge $\mathcal{R}_{C}(Q)$ for context $Q\subseteq\mathcal{K}$ within computational budget $C$ is:
$\mathcal{R}_{C}(Q)=\{k\in\mathcal{K}:\{k\}\not\perp_{C}Q\}.$ (3)
We can now introduce the concept of semantic information measure to describe a general process of reasoning. That special mapping measures information available in any subset of knowledge and extracts its meaning within available computational budget.
@@
The semantic information measure is a bounded mapping
$\mu_{C}:\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathbb{R}_{+}$ (4)
parameterized by computational budget $C$ satisfying:
$\mu_{C}(\emptyset|Q)=0\text{ for all }Q\subseteq\mathcal{K}$ (null set)
$\mu_{C}(A|Q)\geq 0\text{ with }\mu_{C}(A|Q)>0\text{ iff }A\cap\mathcal{R}_{C}(Q)\neq\emptyset$ (contextual non-negativity)
$\mu_{C}(A|Q)\leq\mu_{C}(B|Q)\text{ if }A\subseteq B$ (monotonicity)
$\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)\text{ with equality for }A\perp_{C,Q}B$ (subadditivity)
$\mu_{C}(A|Q)\leq\mu_{D}(A|Q)\text{ if }C<D$ (insight monotonicity)
$\sup\{\mu_{C}(A|Q):A\subseteq\mathcal{K}\}<\infty$ (boundedness)
where $A\perp_{C,Q}B$ means $A\perp_{C}B$ within $\mathcal{R}_{C}(Q)$.
Knowledge that cannot resolve any uncertainty or knowledge irrelevant to a query provides no information. However, if there is uncertainty resolving information within our cognitive reach, we can use it and perform reasoning organizing that information into available knowledge. That is why we assume information content is additive for independent knowledge sets, i.e., when no knowledge in $A$ can be derived from $B$ within computational budget $C$, and vice versa. With redundant information in $A\cup B$ , we admit potential loss of information that may be caused by aggregation. Finally, with increasing computational budget we can extract more knowledge (ordered information) from the same information subset.
Justification for the Axioms of the Semantic Information Measure
The choice of conditional subadditivity for computationally independent sets, i.e., demanding $\mu_{C}(A\cup B|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q)$ with equality for $A\perp_{C}B$, is necessary to describe LLM inference as an auction of ideas. We need to isolate knowledge encoded in distinct neural circuits, attention heads, or activation patterns. The concept of independence is the formal translation of this informational separation into a computational context. It states that the knowledge in idea $A$ cannot be used to derive the knowledge in idea $B$ within a given computational budget $C$, and vice-versa.
@@
However, there is also another form of synergy that we have already included in our model. To see what it may be, we should understand how context $Q$ may affect the semantic measure of any give subset $A$ of knowledge. The semantic information measure can exhibit superadditivity in the context argument.
To see that, suppose $Q_{1}=\{P\}$ and $Q_{2}=\{P\to R\}$, and consider $A=\{P\}$. If computational budget $C=1$ allows one step and we have $f\in\mathcal{F}_{C}$ representing modus ponens, then
$R\cap f(Q_{1})=\emptyset$ $\text{ and }f(R)\cap Q_{1}=\emptyset,$
$R\cap f(Q_{2})=\emptyset$ $\text{ and }f(R)\cap Q_{2}=\emptyset,$
$R\cap f(Q_{1}\cup Q_{2})\neq\emptyset$ $\text{ and }f(R)\cap(Q_{1}\cup Q_{2})\neq\emptyset.$
Therefore, by the properties of semantic information measure:
$\mu_{C}(A|Q_{1}\cup Q_{2})>\mu_{C}(A|Q_{1})+\mu_{C}(A|Q_{2})=0.$
Integrating different pieces of context may synergistically unlock reasoning paths not available individually.
Another important consequence of the assumptions we have made is that we permit structures where $\mu_{C}$ is not monotonic in context. Consider, for the sake of example, $\mu_{C}(A|Q)=|A\setminus Q|$ and $\mathcal{R}_{C}(Q)=\mathcal{K}\setminus Q$. It is rather easy to see that $\mu_{C}$ satisfies the axioms of the semantic information measure. Let us then see what happens if $A=\{k\}$, $Q_{1}=\emptyset$ and $Q_{2}=\{k\}$. We have:
$Q_{1}\subseteq Q_{2}\quad\text{and}\quad\mu_{C}(A|Q_{1})=|\{k\}\setminus\emptyset|=1>0=|\{k\}\setminus\{k\}|=\mu_{C}(A|Q_{2}).$
Increasing the context may reduce the information contributions. If the extended context $(Q_{2})$ provides an alternative explanation for phenomena related to original context $(Q_{1})$, then marginal contribution of available knowledge $(A)$ may decrease.
The axiomatic choices we have made are most applicable to the idealized setting of Theorem 6. Later, in Theorem 8 and 9, we relax them by moving the source of interaction from the measure itself to the aggregation function in actual transformers.
@@ 4.1.1 Limits of Knowledge and Reasoning
For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ there exists
$\mu_{\infty}(A|Q)=\lim_{C\to\infty}\mu_{C}(A|Q).$ (5)
@@
Since $\mu_{C}(A|Q)\leq\mu_{D}(A|Q)$ when $C<D$ and $\mu_{C}$ is bounded on finite $\mathcal{K}$, the monotonic increasing and bounded sequence $\{\mu_{C}(A|Q)\}_{C}$ converges with $C\to\infty$ for any subset of knowledge space. ∎
@@
Let $\mathcal{F}_{C}$ be a complete set of all computational transformations with cost (complexity) lower than or equal to $C$, such that $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ when $C<D$ and such that for any $f,g\in\mathcal{F}_{\infty}=\bigcup_{C}\mathcal{F}_{C}$ there exists $h\in\mathcal{F}_{\infty}$ for which $g(A\cup f(A))=h(A)$. Also, we call $f$ monotone when it is monotone with respect to set inclusion.
The emergence operator is a mapping $\mathcal{E}_{C}\colon\mathcal{P}(\mathcal{K})\times\mathcal{P}(\mathcal{K})\rightarrow\mathcal{P}(\mathcal{K})$ defined as follows:
$\mathcal{E}_{C}(A|Q)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{C}\}.$ (6)
@@
$\mathcal{E}_{\infty}(A)=A\cup\{k\in\mathcal{K}\mid k=f(A)\in\mathcal{R}_{C}(Q)\ \text{for some monotone}\ f\in\mathcal{F}_{\infty}\}.$ (7)
@@
Philosophically, the emergence operator describes reasoning as explicitation, i.e., as the process of making latent implications manifest. When $A\subseteq\mathcal{E}_{C}(A|Q)$, the discovered additional elements represent knowledge that was implicit in $A$ but required computational work $C$ to become accessible. That aligns with the pragmatist or constructivist view that understanding involves active construction rather than passive reception.
When $\mathcal{E}_{C}(A|Q)=A\cup B$, we say that $B$ emerges from reasoning based on $A$. Then $A$ and $B$ are computationally dependent and we write $A\not\perp_{C}B$. In such a case, it follows from the properties of $\mu_{C}$ that we can gain useful information. Whenever $A\subseteq A\cup B$, we have
$\mu_{C}(A|Q)\leq\mu_{C}(A\cup B|Q)=\mu_{C}(\mathcal{E}_{C}(A|Q)|Q)\leq\mu_{C}(A|Q)+\mu_{C}(B|Q).$ (9)
The emergence operator is monotonic, which means the reasoning process it describes is also monotonic. Whenever $A\subseteq B$, we must have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{C}(B|Q)$. Also, we preserve insight monotonicity, because by definition with $\mathcal{F}_{C}\subseteq\mathcal{F}_{D}$ for $C<D$, we have $\mathcal{E}_{C}(A|Q)\subseteq\mathcal{E}_{D}(A|Q)$. By design, it also guarantees existence of a fixed point to which the composition of reasoning rules converges.
@@
For any $A\subseteq\mathcal{K}$ and budget $C>0$, there exists a fixed point of the emergence operator $\mathcal{E}_{C}$:
$A^{*}_{C}=\bigcup_{n\geq 0}^{|\mathcal{K}|}\mathcal{E}_{C}^{(n)}(A|Q)=\mathcal{E}_{C}(A^{*}_{C}|Q).$ (10)
@@
$\mu_{\infty}(A|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(A|Q)|.$ (11)
@@
$\mu_{\infty}(\emptyset|Q)=\log|\mathcal{K}|\cdot|\bigcup_{n=0}^{|\mathcal{K}|}\mathcal{E}^{(n)}_{\infty}(\emptyset|Q)|=\log|\mathcal{K}|\cdot|\emptyset|=0.$
If $A\subseteq B$, then $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(B|Q)$ implies $\mu_{\infty}(A|Q)\leq\mu_{\infty}(B|Q)$. When $A\perp_{\infty}B$, no element of $A^{*}_{\infty}$ can be derived from $B$ (and vice versa). Since $A\cap B=\emptyset$, we conclude $A^{*}_{\infty}\cap B^{*}_{\infty}=\emptyset$ and $|(A\cup B)^{*}_{\infty}|=|A^{*}_{\infty}|+|B^{*}_{\infty}|$. That means $\mu_{\infty}(A\cup B|Q)=\mu_{\infty}(A|Q)+\mu_{\infty}(B|Q)$ for computationally independent sets. That shows the proposed formula satisfies all required properties of semantic information measure.
@@
We say that $X\subseteq\mathcal{K}$ and $Y\subseteq\mathcal{K}$ are equivalent in context $Q\subseteq\mathcal{K}$, i.e., $X\equiv Y$, if and only if $X\subseteq\mathcal{E}_{\infty}(Y|Q)$ and $Y\subseteq\mathcal{E}_{\infty}(X|Q)$.
Notice that the equivalence of knowledge does not mean $A$ is a fixed point, i.e., it is not true that $A=\mathcal{E}_{\infty}(A|Q)$ when $A\equiv\mathcal{E}_{\infty}(A|Q)$. If we have $A=\mathcal{E}_{\infty}(A|Q)$, then nothing more can be derived from $A$ in the context of $Q$ by any reasoning rule from $A$. That statement is very strong and is not true when $A\equiv\mathcal{E}_{\infty}(A|Q)$. Suppose $A=\{x\}$ and $f\in\mathcal{F}_{\infty}$ allows for getting $y=f(x)\in\mathcal{R}_{C}(Q)$. Then $\mathcal{E}_{\infty}(A|Q)=\{x,y\}\supset\{x\}=A\equiv\mathcal{E}_{\infty}(A|Q)$.
@@
For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$ we have:
$\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q).$ (12)
@@
By definition $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$. To show $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$, we must find $h\in\mathcal{F}_{\infty}$ such that $h(A|Q)=g(A\cup\bigcup f_{i}(A|Q))$ for $g,f_{i}\in\mathcal{F}_{\infty}$. But since we can always find such $h\in\mathcal{F}_{\infty}$, we conclude that any context-relevant conclusion $g(\mathcal{E}_{\infty}(A|Q))$ derived from $\mathcal{E}_{\infty}(A|Q)$ by a reasoning rule $g\in\mathcal{F}_{\infty}$ can be also derived directly from $A$ as $h(A)=g(A\cup\bigcup f_{i}(A))$. Therefore, by definition of emergence operator we have $\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))\subseteq\mathcal{E}_{\infty}(A|Q)$ as required. ∎
@@
For any $A\subseteq\mathcal{K}$ and context $Q\subseteq\mathcal{K}$, if the semantic information is constant on equivalent sets of knowledge, i.e., $\mu_{\infty}(X|Q)=\mu_{\infty}(Y|Q)$ whenever $X\equiv Y$, then:
$\mu_{\infty}(\mathcal{E}_{\infty}(A|Q)|Q)=\mu_{\infty}(A|Q).$ (13)
@@
We must show equivalence of the knowledge encoded in $A$ and $\mathcal{E}_{\infty}(A|Q)$, i.e., we demand $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$. But, by Theorem 3, we see that $A\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ and $\mathcal{E}_{\infty}(A|Q)\subseteq\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))=\mathcal{E}_{\infty}(A|Q)$ as requested. Since $A\equiv\mathcal{E}_{\infty}(A|Q)$ and $\mu_{\infty}$ is constant on equivalent sets of knowledge by assumption, we conclude that $\mathcal{E}_{\infty}(A|Q)=\mathcal{E}_{\infty}(\mathcal{E}_{\infty}(A|Q))$. ∎
Justification for the Axiom of Unlimited Reasoning with Emergence Operator
That axiom is essential for the proof of idempotence of unlimited reasoning in Theorem 3 Theorem 4. These are critical philosophical and mathematical results necessary for our study. With them we argue that the information gain observed in practical, constrained inference is an artifact of computational bounds, not the magical creation of knowledge ex nihilo. We acknowledge that depends on the assertion that any sequential reasoning process can be compiled into a single computational step. Therefore, the philosophical claim that information gain is an artifact of computational bounds rests on the acceptance of this structure of reasoning.
4.1.2 The Uncertainty Resolution Paradox and its Resolution
When LLM combines knowledge elements to fabricate facts or produce emergent insights in a given fixed context, it seems that
$\mu_{C}(A\cup B|Q)>\mu_{C}(A|Q)+\mu_{C}(B|Q),$
so novel insights may come from existing data. That is exactly the creativity behind the hallucination phenomenon we are struggling to control. So why should we assume subadditivity?